OpenStudy (anonymous):

Find c and d so that (5-7i)(c+di)=1

OpenStudy (anonymous):

When I FOIL it out, I get 5c+5di-7ci-7di^2=1 But I don't know where to go from there...

OpenStudy (anonymous):

i^2 = -1

OpenStudy (anonymous):

i^2 means -1 and then equate the real and imaginary parts

OpenStudy (anonymous):

yup..

OpenStudy (anonymous):

So... 5c+5di-7ci+7d=1?

OpenStudy (anonymous):

Yes. Now, write that left side so you separate the real and imaginary parts: (5c + 7d) + (5d - 7c)i = 1 For this equality to be true, 2 things must be true. The real part has to be 1, and the imaginary part 0. So, 5c + 7d = 1 and 5d - 7c = 0 Solve that set of equations and you're done.

OpenStudy (anonymous):

Thank you! My teacher didn't explain it that well so I couldn't understand it...